# Residual gauge theory for quanta of surface plasmons

### v3

@@ Abstract
QED serves as the prototypical model of gauge theories, where gauge symmetry and its spontaneous breaking often dictate physical phenomena. [1, 2] In our construction of quantum states using localized photon waves, a particular gauge symmetry of QED, known as residual gauge symmetry, plays a crucial role. Residual gauge transformation corresponds to a constant shift of the gauge field, and the requirement of invariance under this transformation imposes a constraint on physical states, which can be readily solved. Interestingly, the residual gauge symmetry not only imposes constraints on physical states but also reveals a deep connection to thermodynamic concepts. In particular, the associated constraint equation is intimately related to Joule heating. While QED has been employed in previous studies to investigate surface plasmons, [17, 18, 19, 20, 21] our formulation introduces a novel perspective that has not been explored in the existing literature. For example, we demonstrate that this constraint equation emerges from a unitary transformation of the gauge field, generated by an operator associated with Joule heating. The Joule-heating operator represents the fundamental channel through which electromagnetic energy is transferred into the electronic degrees of freedom. This irreversible energy exchange, even if weak, is intrinsic to the existence of plasmon quanta. A perfectly lossless plasmon with an infinite lifetime can be defined theoretically, but such a mode would be completely isolated, contradicting the very coupling that gives rise to it. This suggests a fundamental connection between residual gauge symmetry and Joule heating (or thermodynamics). In a periodic space, a residual gauge transformation corresponds to the Aharonov-Bohm effect, highlighting the physical reality of the gauge field. [22, 23] Therefore, our findings indicate an unexpected link between spatial topology and the physics of surface plasmon quanta, mediated by residual gauge symmetry.
@@ I.2 Boundary condition
Our formulation can be naturally understood within the action principle, starting from the standard QED Lagrangian, ${\cal L}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}+A_{\nu}j^{\nu}$, where $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$[1] We employ the identity $\frac{1}{4}F_{\mu\nu}F^{\mu\nu}=\frac{1}{2}A_{\nu}\partial_{\mu}F^{\mu\nu}-\frac{1}{2}\partial_{\mu}(F^{\mu\nu}A_{\nu})$. In the presence of a 2DEG confined to the plane $z=0$, the total derivative term cannot be discarded, because it does not vanish ($F^{\mu\nu}\neq 0$ and $A_{\nu}\neq 0$) and gives a nontrivial contribution at the boundary. To ensure that the Maxwell equations hold everywhere and that the variational principle is well posed in the presence of a boundary, one must add the explicit boundary counterterm ${\cal L}_{bc}=\partial_{\mu}(F^{\mu\nu}A_{\nu})$. This boundary counterterm is not introduced merely to cancel the total derivative, but to ensure the correct boundary conditions: [40] in the bulk ($z\neq 0$), arbitrary variations of $A_{\nu}$ are allowed, whereas on the boundary, only residual gauge variations ($\delta F^{\mu\nu}=0$ but $\delta A_{\nu}\neq 0$) are physically admissible. With this counterterm, the total Lagrangian becomes ${\cal L}+{\cal L}_{bc}=\frac{1}{2}A_{\nu}\partial_{\mu}F^{\mu\nu}+\frac{1}{2}\partial_{\mu}(F^{\mu\nu}A_{\nu})+A_{\nu}j^{\nu}$. This allows the boundary part of the “bulk” Lagrangian to be written as $\frac{1}{2}A_{\nu}(\partial_{\mu}F^{\mu\nu}+j^{\nu})$. The corresponding action is an extremum under variations of $A_{\nu}$ that satisfy the Maxwell equations $\partial_{\mu}F^{\mu\nu}+j^{\nu}=0$. The action remains stationary under variations of $A_{\nu}$ at $z=0$, i.e., residual gauge variations, because they do not alter $\partial_{\mu}F^{\mu\nu}+j^{\nu}=0$. In the true bulk ($z\neq 0$), we obtain the free-space Maxwell equations $\partial_{\mu}F^{\mu\nu}=0$ by a general variation with respect to $A_{\nu}$, because $j^{\nu}=0$ there. The remaining half of the interaction term combines with the remaining half of the boundary counterterm to form the “boundary” Lagrangian: $\frac{1}{2}{\cal L}_{bc}+\frac{1}{2}A_{\nu}j^{\nu}$. When integrated across an infinitesimal region around $z=0$, this becomes $\frac{1}{2}A_{\nu}([F^{z\nu}]^{0_{+}}_{0_{-}}+J^{\nu})$. The surface integral of this expression leads to the local Hamiltonian given in Eq. (18). This also provides another explanation for why the derivative term $\dot{E}_{i}A_{i}$ at $z=0$ can be neglected. In this derivation, the role of the boundary counterterm is essential for maintaining consistency of the variational procedure and for making the emergence of the constraint transparent, although its interpretation as a Hamiltonian rather than an energy functional remains subtle.
@@ Acknowledgments
I am deeply grateful to Prof. Y. Tokura, whose constant encouragement helped me stay motivated during this research.

### v4

@@ Abstract
QED serves as the prototypical model of gauge theories, where gauge symmetry and its spontaneous breaking often dictate physical phenomena. [1, 2] While QED has been employed in previous studies to investigate surface plasmons, [17, 18, 19, 20, 21] the present formulation provides a framework in which residual gauge symmetry and boundary counterterms are treated explicitly. Residual gauge transformation corresponds to a constant shift of the gauge field, and the requirement of invariance under this transformation imposes a constraint on physical states, which can be readily solved. Interestingly, the residual gauge symmetry not only imposes constraints on physical states but also reveals a deep connection to thermodynamic concepts. For example, we demonstrate that this constraint equation emerges from a unitary transformation of the gauge field, generated by an operator associated with Joule heating. The Joule-heating operator represents the fundamental channel through which electromagnetic energy is transferred into the electronic degrees of freedom. This irreversible energy exchange, even if weak, is intrinsic to the existence of plasmon quanta. A perfectly lossless plasmon with an infinite lifetime can be defined theoretically, but such a mode would be completely isolated, contradicting the very coupling that gives rise to it. This suggests a fundamental connection between residual gauge symmetry and Joule heating (or thermodynamics). In a periodic space, a residual gauge transformation corresponds to the Aharonov-Bohm effect, highlighting the physical reality of the gauge field. [22, 23] Therefore, our findings indicate an unexpected link between spatial topology and the physics of surface plasmon quanta, mediated by residual gauge symmetry.
@@ I.2 Boundary condition
[…]
@@ Acknowledgments
I am deeply grateful to Prof. Y. Tokura for helpful discussions and for his encouragement of this line of research.